### v1

# Improved FPT Approximation Algorithms for TSP

Abstract
TSP is a classic and extensively studied problem with numerous real-world applications in artificial intelligence and operations research. It is well-known that TSP admits a constant approximation ratio on metric graphs but becomes NP-hard to approximate within any computable function $f(n)$ on general graphs. This disparity highlights a significant gap between the results on metric graphs and general graphs. Recent research has introduced some parameters to measure the “distance” of general graphs from being metric and explored FPT approximation algorithms parameterized by these parameters. Two commonly studied parameters are $p$, the number of vertices in triangles violating the triangle inequality, and $q$, the minimum number of vertices whose removal results in a metric graph. In this paper, we present improved FPT approximation algorithms with respect to these two parameters. For $p$, we propose an FPT algorithm with a 1.5-approximation ratio, improving upon the previous ratio of 2.5. For $q$, we significantly enhance the approximation ratio from 11 to 3, advancing the state of the art in both cases.
1 Introduction
The Traveling Salesman Problem (TSP) is a classic optimization problem with numerous real-world applications Applegate (2006), including logistics, genetics, manufacturing, and telecommunications. Given an edge-weighted complete graph $G=(V,E,w)$ with $n$ vertices in $V$ and a non-negative weight function $w:E\to\mathbb{R}_{\geq 0}$ on edges, the goal of TSP is to find an optimal (i.e., minimum-weight) TSP tour (i.e., Hamiltonian cycle) that visits each vertex exactly once. Due to its NP-hardness Garey and Johnson (1979), TSP has been extensively studied in the context of approximation algorithms Saller et al. (2023); Traub and Vygen (2025).
For a minimization problem, an algorithm is called a $\rho$-approximation algorithm if, for any instance $\mathcal{I}$, it computes a solution with an objective value that is at most $\rho$ times the optimal value in polynomial time $poly(|\mathcal{I}|)$. Here, $\rho$ is called the approximation ratio. Moreover, with some parameter $k$, an algorithm is called an FPT $\rho$-approximation algorithm if, for any instance $\mathcal{I}$, it runs in $\mathcal{O}(f(k))\cdot poly(|\mathcal{I}|)$ time with an approximation ratio of $\rho$, where $f(k)$ is a computable function of $k$. Both FPT algorithms and FPT approximation algorithms are particularly useful in practice when the parameter $k$ is small Cygan et al. (2015).
While TSP on general graphs is NP-hard to approximate within any computable function $f(n)$ Sahni and Gonzalez (1976), TSP on metric graphs (metric TSP), where the weight function satisfies the triangle inequality, admits a constant approximation ratio. Notably, Christofides Christofides (2022) and Serdyukov Serdyukov (1978) independently proposed the famous $1.5$-approximation algorithm for metric TSP. Currently, Karlin et al. Karlin et al. (2021, 2023) have improved the approximation ratio to $1.5-$ $10^{-36}$. Karpinski et al. Karpinski et al. (2015) showed that metric TSP is NP-hard to approximate within a ratio of $123/122$.
Given the significant gap between the results on general graphs and metric graphs, many approximation algorithms have been developed for graphs that are “nearly” metric, aiming to measure the “distance” from approximability. In the literature, two main research directions are commonly explored. The first direction relaxes the condition on metric functions, where the weight function satisfies the $\tau$-triangle inequality: $\tau\cdot(w(a,b)+w(b,c))\geq w(a,c)$ for all $a,b,c\in V$. The second direction relaxes the scope of metric functions, where the weight function partially satisfies the triangle inequality, i.e., only a subset of all triangles is required to satisfy it.
For the first direction, Andreae and Bandelt Andreae and Bandelt (1995) gave a $(3\tau^{2}+\tau)/2$-approximation algorithm for TSP. Bender and Chekuri Bender and Chekuri (2000) improved this ratio to $4\tau$ when $\tau>7/3$, and Mömke Mömke (2015) further improved the ratio to $43(\tau^{2}+\tau)/$ 4 for small values of $\tau$. Additional improvements Andreae (2001); Böckenhauer et al. (2002) can be found in a survey Klasing and Mömke (2018).
For the second direction, Mohan Mohan et al. (2017) proposed a $3.5$-approximation algorithm for TSP on graphs $G=(V_{1}\cup V_{2},$ $E,w)$, where both induced subgraphs $G[V_{1}]$ and $G[V_{2}]$ are required to be metric graphs. Sumita et al. Sumita et al. (2017) proposed a $3$-approximation algorithm for Subgroup Path Planning, where the vertex set $V$ is partitioned into $k$ disjoint groups $V_{1},...,V_{k}$. The goal is to find an optimal TSP tour visiting the vertices in each group $V_{i}$ consecutively, where the weight function violates $w(a,b)+w(b,c)\geq w(a,c)$ only if both $a$ and $c$ belong to the same group, while $b$ belongs to a different group. Moreover, if it holds that $|V_{i}|=2$ for all $i$, the problem reduces to Subpath Planning Safilian et al. (2016), which admits a $1.5$-approximation algorithm Sumita et al. (2017). Both problems have various applications in AI, such as polishing robots and electronic printing Safilian et al. (2016); Sumita et al. (2017).
Subsequently, Zhou et al. Zhou et al. (2022, 2024) introduced two natural parameters: the number of vertices $p$ in violating triangles that do not satisfy the triangle inequality, and the minimum number of violating vertices $q$ whose removal results in a metric graph. Note that $q\leq p$. Moreover, the parameter $p$ can be computed in polynomial time, while $q$ can be computed in $\mathcal{O}(3^{q}n^{3})$ time using the search-tree method Zhou et al. (2022). For TSP parameterized by $p$, they proposed an FPT $3$-approximation algorithm with a running time of $2^{\mathcal{O}({p\log p})}\cdot n^{\mathcal{O}(1)}$. For TSP parameterized by $q$, they proposed a $3$-approximation algorithm with a running time of $n^{\mathcal{O}(q+1)}$, and an FPT $11$-approximation algorithm with a running time of $2^{\mathcal{O}({q\log q})}\cdot n^{\mathcal{O}(1)}$. Recently, Bampis et al. Bampis et al. (2024) improved the FPT approximation ratio from $3$ to $2.5$ for the parameter $p$, while maintaining the running time of $2^{\mathcal{O}({p\log p})}\cdot n^{\mathcal{O}(1)}$. Our research follows this line to study FPT approximation algorithms for TSP with respect to these two parameters.
1.1 Our results
In this paper, we present improved FPT approximation algorithms for TSP with respect to the parameters $p$ and $q$, addressing the question raised in Bampis et al. (2024) regarding the potential for further improvement in the approximation ratio. Let $\alpha$ denote the approximation ratio for metric TSP. Currently, the best-known value of $\alpha$ is $\alpha=1.5-10^{-36}$ Karlin et al. (2023). Our results are summarized as follows.
For TSP parameterized by $p$, we first introduce a simple $(\alpha+1)$-approximation algorithm with a running time of $2^{\mathcal{O}(p)}+n^{\mathcal{O}(1)}$. This improves upon the previous 2.5-approximation algorithm, which has a worse running time of $2^{\mathcal{O}(p\log p)}\cdot n^{\mathcal{O}(1)}$ Bampis et al. (2024). Additionally, we propose a more sophisticated $1.5$-approximation algorithm, though it comes with a running time of $2^{\mathcal{O}(p\log p)}\cdot n^{\mathcal{O}(1)}$.
For TSP parameterized by $q$, we present an FPT $3$-approximation algorithm, significantly improving upon the previous FPT $11$-approximation algorithm Zhou et al. (2024). Both algorithms run in $2^{\mathcal{O}(q\log q)}\cdot n^{\mathcal{O}(1)}$ time. Our result also matches the approximation ratio of the earlier algorithm with a running time of $n^{\mathcal{O}(q+1)}$ Zhou et al. (2022).
A summary of the previous and our results is provided in Table 1. Our improvements arise from the application of novel techniques for handling non-metric graphs, which may also prove valuable for tackling other related problems. Detailed comparisons between our algorithms and previous approaches will be provided later.
Due to space constraints, the proofs of the lemmas and theorems marked with $\clubsuit$ have been moved to the appendix.
Para. Approx. Runtime Reference
$p$ $3$ $2^{\mathcal{O}({p\log p})}\cdot n^{\mathcal{O}(1)}$ Zhou et al. Zhou et al. (2022)
$2.5$ $2^{\mathcal{O}({p\log p})}\cdot n^{\mathcal{O}(1)}$ Bampis et al. Bampis et al. (2024)
$\alpha+1$ ${2^{\mathcal{O}(p)}+n^{\mathcal{O}(1)}}$ This Paper
1.51.5{1.5}1.5 $2^{\mathcal{O}({p\log p})}\cdot n^{\mathcal{O}(1)}$
$q$ $3$ $n^{\mathcal{O}(q+1)}$ Zhou et al. Zhou et al. (2022)
$11$ $2^{\mathcal{O}({q\log q})}\cdot n^{\mathcal{O}(1)}$ Zhou et al. Zhou et al. (2024)
33{3}3 $2^{\mathcal{O}({q\log q})}\cdot n^{\mathcal{O}(1)}$ This Paper
Table 1: A summary of previous results and our results on approximation algorithms for TSP parameterized by $p$ and $q$.
2 Preliminaries
Let $G=(V,E,w)$ denote the input complete graph, where there are $n$ vertices in $V$ and a weight function $w:E\to\mathbb{R}_{\geq 0}$. For any $E^{\prime}\subseteq E$, define $w(E^{\prime})\coloneqq\sum_{e\in E^{\prime}}w(e)$. $G$ is called a metric graph if its weight function satisfies the symmetry and triangle inequality properties, i.e., $w(a,b)=w(b,a)\leq w(a,c)+w(c,b)$ for all $a,b,c\in V$. A triangle $\Delta(a,b,c)$ in $G$ is called a violating triangle if it violates the triangle inequality. For any subgraph $G^{\prime}$ of $G$, let $V(G^{\prime})$ (resp., $E(G^{\prime})$) denote the vertex (resp., edge) set. For any $V^{\prime}\subseteq V(G^{\prime})$, let $G^{\prime}[V^{\prime}]$ denote the subgraph of $G^{\prime}$ induced by $V^{\prime}$. For any $V^{\prime}\subseteq V$, it is called a violating set of $G$ if $G[V\setminus V^{\prime}]$ is a metric graph, and the vertices in $V^{\prime}$ are called violating vertices. A spanning $t$-forest $\mathcal{F}$ in $G$ is a set of vertex-disjoint trees with $|\mathcal{F}|=t$ and $V(\mathcal{F})=V$. Note that an isolated vertex is also a tree. A matching $\mathcal{M}$ in $G$ is a set of vertex-disjoint edges with $V(\mathcal{M})=V$. For any integer $t$, let $[t]\coloneqq\{1,...,t\}$.
For any (multi-)graph $G^{\prime}=(V^{\prime},E^{\prime},w^{\prime})$, the degree of a vertex is the number of edges incident to it. Let $Odd(G^{\prime})$ denote the set of odd-degree vertices in $V^{\prime}$. Then, $G^{\prime}$ is called an Eulerian graph if it is a connected graph and every vertex in $V^{\prime}$ has an even degree. In $G^{\prime}$, a tour $T=\overline{v_{1}v_{2}...v_{s}}$ is a sequence of vertices, where a vertex may appear multiple times. $T$ consists of $s$ edges in $E^{\prime}(T)\coloneqq\{v_{1}v_{2},v_{2}v_{3},...,v_{s}v_{1}\}$, and let $w^{\prime}(T)\coloneqq w^{\prime}(E^{\prime}(T))$. Moreover,
  • $T$ is called a simple tour if it contains no repeated vertices, and it is called a TSP tour if it further satisfies that $|E^{\prime}(T)|=|V^{\prime}|$;
  • $T$ is called an Eulerian tour if $E^{\prime}(T)$ is exactly the set of all edges in $G^{\prime}$. If $G^{\prime}$ is an Eulerian graph, an Eulerian tour can be computed in $\mathcal{O}(|E^{\prime}|)$ time Cormen et al. (2022). We may use $G^{\prime}$ to denote an Eulerian tour in $G^{\prime}$.
Given a tour $T=\overline{...xyz...}$, we can obtain a new tour $\overline{...xz...}$ by skipping the vertex $y$. This operation is called shortcutting the vertex $y$ or the edges $xy$ and $yz$, or taking a shortcut on $T$. Given a simple tour $\overline{v_{1}v_{2}...v_{s}}$ in $G^{\prime}$, we use $A=v_{1}v_{2}...v_{s}$ to denote a $s$-path, which consists of $s-1$ edges in $E^{\prime}(A)\coloneqq\{v_{1}v_{2},v_{2}v_{3},...,v_{s-1}v_{s}\}$. Note that a $1$-path is a single vertex. A $|V^{\prime}|$-path in $G^{\prime}$ is also called a TSP path, and we also refer to a path as a chain.
In the following, we fix an optimal TSP tour $T^{*}$ in $G$, and denote the weight of $T^{*}$ by OPT.
3 TSP Parameterized by the Number of Vertices in Violating Triangles
We first define some notations. A vertex is called a bad vertex if it appears in a violating triangle, and a good vertex otherwise. Let $V=V_{b}\cup V_{g}$, where $V_{b}$ (resp., $V_{g}$) denotes the set of bad (resp., good) vertices in $V$. Note that $|V_{b}|=p$. Assume w.l.o.g. that $\min\{|V_{g}|,|V_{b}|\}\geq 3$. Otherwise, if $|V_{g}|<3$, the problem can be solved optimally in $2^{\mathcal{O}(p)}$ time using the dynamic programming (DP) method Bellman (1962); Held and Karp (1962); if $|V_{b}|<3$, the number of violating triangles is 0, i.e., $|V_{b}|=0$, so the problem reduces to metric TSP. Thus, when $\min\{|V_{g}|,|V_{b}|\}<3$, the problem admits an FPT approximation ratio of at most $\alpha<1.5$ with a running time of $2^{\mathcal{O}(p)}+n^{\mathcal{O}(1)}$.
By definition, we have the following property.
Property 1.
Every triangle containing a good vertex satisfies the triangle inequality.
3.1 The first algorithm
In this subsection, we propose an FPT $(\alpha+1)$-approximation algorithm (ALG.1) with a running time of $2^{\mathcal{O}(p)}+n^{\mathcal{O}(1)}$.
Our ALG.1 is surprisingly simple. First, it arbitrarily selects a good vertex $o\in V_{g}$. Then, it computes an optimal TSP tour $T_{b}$ in $G[V_{b}\cup\{o\}]$ using the DP method Bellman (1962); Held and Karp (1962). Next, it computes an $\alpha$-approximate TSP tour $T_{g}$ in $G[V_{g}]$ by applying an $\alpha$-approximation algorithm for metric TSP. Finally, it constructs a TSP tour $T_{1}$ in $G$ by taking a shortcut on $T_{b}\cup T_{g}$. ALG.1 is described in Alg.1.
Note that the TSP tour $T_{b}$ in ALG.1 contains one good vertex. This trick enables us to take a shortcut on $T_{b}\cup T_{g}$ to get a TSP tour in $G$ without increasing the weight (see Theorem 3).
Algorithm 1 ALG.1
Input: An instance $G=(V=V_{g}\cup V_{b},E,w)$.
Output: A feasible solution.
1: Arbitrarily select a good vertex $o\in V_{g}$.
2: Compute an optimal TSP tour $T_{b}$ in $G[V_{b}\cup\{o\}]$ using the DP method Bellman (1962); Held and Karp (1962).
3: Compute an $\alpha$-approximate TSP tour $T_{g}$ in $G[V_{g}]$ by applying an $\alpha$-approximation algorithm for metric TSP.
4: Take a shortcut on $T_{b}\cup T_{g}$ to obtain a TSP tour $T_{1}$ in $G$.
Lemma 2.
It holds that $\mbox{OPT}\geq w(T^{*}_{b})$ and $\mbox{OPT}\geq w(T^{*}_{g})$, where $T^{*}_{b}$ (resp., $T^{*}_{g}$) denotes an optimal TSP tour in $G[V_{b}\cup\{o\}]$ (resp., $G[V_{g}]$).
Proof.
First, we prove $\mbox{OPT}\geq w(T^{*}_{b})$. By Property 1, every triangle containing a good vertex satisfies the triangle inequality. So, given the optimal TSP tour $T^{*}$ in $G$, we can directly shortcut all good vertices in $V_{g}\setminus\{o\}$ to get a TSP tour $T^{\prime}$ in $G[V_{b}\cup\{o\}]$, such that $\mbox{OPT}\geq w(T^{\prime})$. Since $T^{*}_{b}$ is an optimal TSP tour in $G[V_{b}\cup\{o\}]$, we must have $\mbox{OPT}\geq w(T^{\prime})\geq w(T^{*}_{b})$.
Next, we prove $\mbox{OPT}\geq w(T^{*}_{g})$. Consider the optimal TSP tour $T^{*}$ in $G$. Recall that by our assumption, $\min\{|V_{g}|,|V_{b}|\}\geq 3$. Thus, there must be a bad vertex that is incident to a good vertex in $T^{*}$. By Property 1, we can use the good vertex to shortcut the bad vertex without increasing the weight. Thus, by shortcutting all bad vertices one by one, we obtain a TSP tour $T^{\prime\prime}$ in $G[V_{g}]$ with $\mbox{OPT}\geq w(T^{\prime\prime})$. Similarly, we have $\mbox{OPT}\geq w(T^{\prime\prime})\geq w(T^{*}_{g})$. ∎
Theorem 3.
For TSP parameterized by $p$, ALG.1 is an FPT $(\alpha+1)$-approximation algorithm with runtime $2^{\mathcal{O}(p)}+n^{\mathcal{O}(1)}$.
Proof.
First, we analyze the solution quality of ALG.1. Since $T_{1}$ is obtained by taking a shortcut on $T_{b}\cup T_{g}$, we assume w.l.o.g. that $T_{1}=\overline{ou_{1}...u_{p}v_{1}...v_{n-p-1}}$, where $T_{b}=\overline{ou_{1}...u_{p}}$ and $T_{g}=\overline{ov_{1}...v_{n-p-1}}$. Since $o$ is a good vertex, the triangle $\Delta(o,v_{1},u_{p})$ satisfies the triangle inequality by Property 1, and then we have $w(u_{p},o)+w(o,v_{1})\geq w(u_{p},v_{1})$. Therefore, we have $w(T_{b})+w(T_{g})\geq w(T_{1})$. Since $T_{b}$ (resp., $T_{g}$) is an optimal (resp., $\alpha$-approximate) TSP tour in $G[V_{b}\cup\{o\}]$ (resp., $G[V_{g}]$), by Lemma 2, we have $(\alpha+1)\cdot\mbox{OPT}\geq w(T_{1})$.
Next, we analyze the running time of ALG.1. It can be verified that Steps 1, 3, and 4 take polynomial time, $n^{\mathcal{O}(1)}$, whereas Step 2 is dominated by computing the optimal TSP tour. Since there are $|V_{b}|+1=p+1$ vertices in $G[V_{b}\cup\{o\}]$, the optimal TSP tour can be computed in $2^{\mathcal{O}(p)}$ time using the DP method Bellman (1962); Held and Karp (1962). Thus, the overall running time of ALG.1 is at most $2^{\mathcal{O}(p)}+n^{\mathcal{O}(1)}$. ∎
By Theorem 3, using the $(1.5-10^{-36})$-approximation algorithm Karlin et al. (2021) for metric TSP, ALG.1 achieves an FPT approximation ratio of $2.5-10^{-36}$, improving upon the best-known algorithm Bampis et al. (2024) in both the approximation ratio and the running time.
3.2 The second algorithm
In this subsection, we propose an FPT $1.5$-approximation algorithm (ALG.2) with a running time of $2^{\mathcal{O}(p\log p)}\cdot n^{\mathcal{O}(1)}$.
The main idea of ALG.2 is as follows. Consider the optimal TSP tour $T^{*}$ in $G$. First, to handle bad vertices, ALG.2 “guesses” the structure of the bad vertices in $T^{*}$ [^1] Throughout this paper, “guess” means that the algorithm runs each possible structure, and returns the best found over all the runs.. Specifically, after deleting all good vertices and the edges incident to good vertices in $T^{*}$, we obtain a set of paths in $G[V_{b}]$, referred to as a set of bad chains $\mathcal{A}$. Since $|V_{b}|=p$, $\mathcal{A}$ can be guessed within $2^{\mathcal{O}(p\log p)}$ loops. Next, ALG.2 aims to augment the bad chains in $\mathcal{A}$ into a TSP tour in $G$. To achieve this, ALG.2 first augments $\mathcal{A}$ into a minimum constrained spanning tree $F_{\mathcal{A}}$ in $G$ such that $E(\mathcal{A})\subseteq E(F_{\mathcal{A}})$, each $i$-degree vertex in $\mathcal{A}$, where $i\in\{0,1,2\}$, is incident to exactly $i$ bad vertices in $F_{\mathcal{A}}$, and each $2$-degree vertex in $\mathcal{A}$ remains a $2$-degree vertex in $F_{\mathcal{A}}$. Then, it finds a set of edges $M_{\mathcal{A}}$ to fix the parity of the degrees of the vertices in $Odd(F_{\mathcal{A}})$, and constructs a TSP tour $T_{2}$ in $G$ by taking shortcuts on $F_{\mathcal{A}}\cup M_{\mathcal{A}}$. ALG.2 is described in Alg.2. An example of ALG.2 can be found in Fig. 1.
Algorithm 2 ALG.2
Input: An instance $G=(V=V_{g}\cup V_{b},E,w)$.
Output: A feasible solution.
1: Guess the set of bad chains $\mathcal{A}$ in $T^{*}$. Denote the set of $i$-degree vertices in $\mathcal{A}$ by $V^{i}_{b}$, where $i\in\{0,1,2\}$.
2: Obtain a complete graph $\widetilde{G}=(V_{\mathcal{A}}\cup V_{g},\widetilde{E},\widetilde{w})$ by contracting each $A\in\mathcal{A}$ into a vertex $u_{A}$, where $V_{\mathcal{A}}$ denotes the set of contracted vertices, and $\widetilde{w}$ is defined as follows: (1) $\widetilde{w}(u_{A},v)=\min\{w(u,v),w(u^{\prime},v)\}$ for any $A=u...u^{\prime}\in\mathcal{A}$ and any $v\in V_{g}$; (2) $\widetilde{w}(v,v^{\prime})=w(v,v^{\prime})$ for any $v,v^{\prime}\in V_{g}$; (3) $\widetilde{w}(u_{A},u_{A^{\prime}})=+\infty$ for any $A,A^{\prime}\in\mathcal{A}$.
3: Find a minimum spanning tree $\widetilde{F}$ in $\widetilde{G}$, which corresponds to a set of edge $F$ in $G$, where $w(F)=\widetilde{w}(\widetilde{F})$.
4: Obtain a constrained spanning tree $F_{\mathcal{A}}=F\cup\mathcal{A}$ in $G$.
5: Obtain a complete graph $G^{\prime}=(V^{\prime}=Odd(F_{\mathcal{A}}),E^{\prime},w^{\prime})$: For any $u,u^{\prime}\in V^{\prime}$, set $w^{\prime}(u,u^{\prime})=w(A)$ if $u...u^{\prime}=A\in\mathcal{A}$, and $w^{\prime}(u,u^{\prime})=w(u,u^{\prime})$ otherwise.
6: Find a minimum matching $M^{\prime}_{\mathcal{A}}$ in $G^{\prime}$, which corresponds to a set of edges $M_{\mathcal{A}}$ in $G$, where $w(M_{\mathcal{A}})=w^{\prime}(M^{\prime}_{\mathcal{A}})$.
7: Obtain a TSP tour $T_{2}$ in $G$ by taking shortcuts on $F_{\mathcal{A}}\cup M_{\mathcal{A}}$ (see Lemma 6).
(a) $F_{\mathcal{A}}=F\cup\mathcal{A}$
(b) $F_{\mathcal{A}}\cup M_{\mathcal{A}}$
(c) $T^{\prime}_{\mathcal{A}}$
(d) $T_{2}$
Figure 1: An illustration of our ALG.2, where each black (resp., white) node represents a bad (resp., good) vertex. In (a), the edges in $E(F)$ (resp., $E(\mathcal{A})$) are represented by red edges (resp., black edges); In (b), the edges in $M_{\mathcal{A}}$ are represented by blue edges; In (c), $\mathcal{T}^{\prime}_{\mathcal{A}}$ is obtained by shortcutting the repeated edges on the paths in $\mathcal{A}\cap\mathcal{A}^{\prime}$ (see Lemma 6), and the new edges are represented by purple edges; In (d), the TSP tour $T_{2}$ is obtained by taking shortcuts on $\mathcal{T}^{\prime}_{\mathcal{A}}$ (see Lemma 6), and the new edges are also represented by purple edges.
Before analyzing ALG.2, we first provide a comparison between our algorithm and previous algorithms.
[…]
We can see that one key difference in ALG.2 is that, instead of finding a spanning $|\mathcal{A}|$-forest in $G[V^{\prime}_{b}\cup V_{g}]$, ALG.2 finds a constrained spanning tree $F_{\mathcal{A}}$ in $G$, which enables us to avoid the need for $E^{\prime}$. However, a challenge arises because $G[Odd(F_{\mathcal{A}})]$ is not a metric graph, so we cannot directly find a minimum matching in $G[Odd(F_{\mathcal{A}})]$ to fix the parity of the degrees of the vertices in $Odd(F_{\mathcal{A}})$. This contrasts with the previous FPT $2.5$-approximation algorithm in Bampis et al. (2024), where $G[Odd(\mathcal{A}\cup E^{\prime}\cup\mathcal{F})]$ has some good properties that allow for the direct computation of a minimum matching. To address this issue, we construct the graph $G^{\prime}$ in Step 5 of ALG.2 and obtain $M_{\mathcal{A}}$ by finding a minimum matching in $G^{\prime}$ to fix the parity of the degrees of the vertices in $Odd(F_{\mathcal{A}})$. Since we are dealing with non-metric graphs, we must carefully take shortcuts on $F_{\mathcal{A}}\cup M_{\mathcal{A}}$ to obtain a TSP tour. Therefore, our ALG.2 uses several new ideas, and we also need to use more techniques to analyze the final approximation ratio.
Next, we analyze the approximation ratio of our ALG.2.
Lemma 4 ($\clubsuit$).
$F_{\mathcal{A}}$ in ALG.2 is a minimum constrained spanning tree in $G$ such that $E(\mathcal{A})\subseteq E(F_{\mathcal{A}})$, each $i$-degree vertex in $\mathcal{A}$, where $i\in\{0,1,2\}$, is incident to exactly $i$ bad vertices in $F_{\mathcal{A}}$, and each $2$-degree vertex in $\mathcal{A}$ remains a $2$-degree vertex in $F_{\mathcal{A}}$. Moreover, it holds that $\mbox{OPT}\geq w(F_{\mathcal{A}})$.
Lemma 5 ($\clubsuit$).
It holds that $\frac{1}{2}\cdot\mbox{OPT}\geq w(M_{\mathcal{A}})$.
Lemma 6.
There is an $\mathcal{O}(n)$-time algorithm to obtain a TSP tour $T_{2}$ in $G$ with weight at most $w(F_{\mathcal{A}})+w(M_{\mathcal{A}})$.
Proof.
Clearly, $F_{\mathcal{A}}\cup M_{\mathcal{A}}$ forms an Eulerian graph $G_{\mathcal{A}}$ with $V(G_{\mathcal{A}})=V$ and $\mathcal{O}(n)$ edges, and therefore, an Eulerian tour $T_{\mathcal{A}}$ can be computed in $\mathcal{O}(n)$ time Cormen et al. (2022).
We observe that
  1. (1)
    the edges between bad vertices in $E(F_{\mathcal{A}})$ are exactly the edges in $E(\mathcal{A})$;
  2. (2)
    the edges between bad vertices in $E(M_{\mathcal{A}})$ form a set of vertex-disjoint paths $\mathcal{A}^{\prime}$ since $M^{\prime}_{\mathcal{A}}$ is a matching in $G^{\prime}$.
Then, by Step 5 of ALG.2, it holds that
  1. (3)
    for each $A\in\mathcal{A}^{\prime}$, either $A\in\mathcal{A}$ or $A$ is edge-disjoint from all paths in $\mathcal{A}$.
Thus, the edges on the paths in $\mathcal{A}\cap\mathcal{A}^{\prime}$ (resp., $\mathcal{A}\setminus\mathcal{A}^{\prime}$) appear exactly twice (resp., once) in $T_{\mathcal{A}}$ (see (b) in Fig.1).
Next, we first show how to shortcut the repeated edges on the paths in $\mathcal{A}\cap\mathcal{A}^{\prime}$. Consider any path $ab...b^{\prime}a^{\prime}\in\mathcal{A}\cap\mathcal{A}^{\prime}$. By (1), (2), and (3), in $V_{b}$, there is only one vertex $b\in V_{b}$ (resp., $b^{\prime}\in V_{b}$) that is incident to $a$ (resp., $a^{\prime}$). Moreover, since $G_{\mathcal{A}}$ is an Eulerian graph, $G_{\mathcal{A}}$ is connected and both $a$ and $a^{\prime}$ have an even degree in $G_{\mathcal{A}}$. So, either $a$ or $a^{\prime}$, say $a$, is incident to at least two good vertices. Thus, we can use one of the good vertices to shortcut the repeated edges on the path $ab...b^{\prime}a^{\prime}$ without increasing the weight by Property 1. Denote the new Eulerian tour by $T^{\prime}_{\mathcal{A}}$. Recall that $Odd(F_{\mathcal{A}})\subseteq V^{0}_{b}\cup V^{1}_{b}\cup V_{g}$ by the proof of Lemma 5, we know that
  1. (4)
    each edge between bad vertices appears once, and each bad vertex is incident to at most two bad vertices (see (c) in Fig.1).
Then, we show how to shortcut the repeated bad vertices in $T^{\prime}_{\mathcal{A}}$. Assume that there is a bad vertex $a$ that appears at least twice in $T^{\prime}_{\mathcal{A}}$. By (4), every appearance of $a$, except for one, is incident to at least one good vertex in $T^{\prime}_{\mathcal{A}}$. Therefore, each such appearance can be shortcut by using an adjacent good vertex without increasing the weight by Property 1.
Last, the repeated good vertices can clearly be shortcut without increasing the weight by Property 1. Therefore, given $F_{\mathcal{A}}\cup M_{\mathcal{A}}$, we can obtain a TSP tour $T_{2}$ in $G$ with weight at most $w(F_{\mathcal{A}})+w(M_{\mathcal{A}})$, and the running time is $\mathcal{O}(n)$. ∎
Theorem 7 ($\clubsuit$).
For TSP parameterized by $p$, ALG.2 has an FPT $1.5$-approximation ratio with runtime $2^{\mathcal{O}(p\log p)}\cdot n^{\mathcal{O}(1)}$.
4 TSP Parameterized by the Number of Violating Vertices
In this section, a vertex is called a bad vertex if it is a violating vertex, and a good vertex otherwise. Let $V=V_{b}\cup V_{g}$, where $V_{b}$ (resp., $V_{g}$) denotes the set of bad (resp., good) vertices in $V$ and $|V_{b}|=q$. Recall that $V_{b}$ is a minimum violating set, which can be computed in $\mathcal{O}(3^{q}n^{3})$ Zhou et al. (2022). We assume w.l.o.g. that $\min\{|V_{g}|,|V_{b}|\}\geq 1$. Then, we define some notations related to the optimal TSP tour $T^{*}$.
In $T^{*}$, all edges between bad vertices form a set of bad chains $\mathcal{A}$, the set of all edges between one bad vertex and one good vertex is denoted by $\mathcal{B}$, where each edge in $\mathcal{B}$ is called a limb, and all edges between good vertices form a set of good chains $\mathcal{R}$. Note that
$\mbox{OPT}=w(\mathcal{A})+w(\mathcal{B})+w(\mathcal{R})\ \ \mbox{and}\ \ |\mathcal{A}|=2|\mathcal{B}|=|\mathcal{R}|.$ (1)
Denote the set of the vertices in $V_{g}$ that are incident to $i$ bad vertices in $T^{*}$ by $V^{i}_{g}$, where $i\in\{0,1,2\}$. The vertices in $V^{0}_{g}$ (resp., $V^{1}_{g}\cup V^{2}_{g}$) are called internal vertices (resp., anchors). An anchor is called a single anchor if it is in $V^{2}_{g}$ and a pair anchor otherwise. Let $V_{a}=V^{1}_{g}\cup V^{2}_{g}$ denote the set of all anchors. We have
$|\mathcal{A}|=2|\mathcal{B}|=|\mathcal{R}|\leq q\quad\mbox{and}\quad|V_{a}|\leq 2q.$ (2)
An illustration of our notations can be found in Fig.2.
Figure 2: An illustration of our notations: the cycle denotes $T^{*}$, each black (resp., white) node denotes a bad (resp., an internal) vertex, and each green node denotes an anchor; $\mathcal{A}=\{x_{1}...x_{4},x_{5}x_{6},x_{7}\}$, =$\mathcal{B}=\{x_{4}y_{1},y_{4}x_{5},x_{6}y_{5},...,y_{8}x_{1}\}$, and =$\mathcal{R}=\{y_{1}...y_{4},y_{5},y_{6}y_{7}y_{8}\}$.
4.1 The algorithm
In this subsection, we propose an FPT $3$-approximation algorithm (ALG.3) with a running time of $2^{\mathcal{O}(q\log q)}\cdot n^{\mathcal{O}(1)}$.
The high-level idea of ALG.3 is as follows. Since a triangle containing one bad vertex and two good vertices may violate the triangle inequality, Property 1 in the previous section does not hold. Thus, only guessing $\mathcal{A}$ may not be enough to obtain a constant FPT approximation ratio. Moreover, since $|V_{g}|$ can be arbitrarily larger than $q$, we cannot guess $\mathcal{B}$ in FPT time. To address these issues, we use an idea in Zhou et al. (2024) to guess a set of limbs $\mathcal{B}^{\prime}$ in FPT time by finding a minimum spanning $|\mathcal{A}|$-forest $\mathcal{F}$ in $G[V_{g}]$ and then guessing a set of anchors in a subset of $V(\mathcal{F})$, where $\mathcal{B}^{\prime}$ is not necessarily equals to $\mathcal{B}$ but it satisfies $w(\mathcal{B}^{\prime})\leq w(\mathcal{B})$. Clearly,
$|\mathcal{F}|\leq q\quad\mbox{and}\quad w(\mathcal{F})\leq w(\mathcal{R}).$ (3)
Then, we aim to augment $\mathcal{A}\cup\mathcal{B}^{\prime}\cup\mathcal{F}$ into an Eulerian graph. The set of good chains $\mathcal{R}$ guarantees the existence of a set of edges $\mathcal{R}^{\prime}$ with $w(\mathcal{R}^{\prime})\leq w(\mathcal{R})$ to augment $\mathcal{A}\cup\mathcal{B}^{\prime}\cup\mathcal{F}$ into a connected graph, and we will prove that $\mathcal{R}^{\prime}$ can be guessed in FPT time. Let $G_{\mathcal{A}}=\mathcal{A}\cup\mathcal{B}^{\prime}\cup\mathcal{R}^{\prime}\cup\mathcal{F}$, which forms a connected graph. Then, we will show that $|Odd(G_{\mathcal{A}})\cap V(F)|$ is even for each $F\in\mathcal{F}$. Therefore, we can find a minimum matching $M_{F}$ in $G[Odd(G_{\mathcal{A}})\cap V(F)]$ for each $F\in\mathcal{F}$ to fix the parity of the degrees of the odd-degree vertices in $G_{\mathcal{A}}$. We will prove $w(M_{F})\leq w(F)$. Let $\mathcal{M}=\bigcup_{F\in\mathcal{F}}M_{F}$. Finally, we obtain an Eulerian graph $G^{\prime}_{\mathcal{A}}=\mathcal{A}\cup\mathcal{B}^{\prime}\cup\mathcal{R}^{\prime}\cup\mathcal{F}\cup\mathcal{M}$, and then we carefully take shortcuts on $G^{\prime}_{\mathcal{A}}$ to obtain a TSP tour without increasing the weight. Thus, by (1), our TSP tour has a weight of at most $w(G^{\prime}_{\mathcal{A}})\leq w(\mathcal{A})+w(\mathcal{B}^{\prime})+w(\mathcal{R}^{\prime})+w(\mathcal{F})+w(\mathcal{M})\leq w(\mathcal{A})+w(\mathcal{B})+w(\mathcal{R})+2w(\mathcal{R})\leq 3\cdot\mbox{OPT}$.
ALG.3 is described in Alg.3, where LIMB, CONNECT, and SHORTCUT are sub-algorithms explained later. An illustration of ALG.3 can be found in Fig. 3.
Algorithm 3 ALG.3
Input: An instance $G=(V=V_{g}\cup V_{b},E,w)$.
Output: A feasible solution.
1: Guess the set of bad chains $\mathcal{A}$ in $T^{*}$.
2: Find a minimum spanning $|\mathcal{A}|$-forest $\mathcal{F}$ in $G[V_{g}]$.
3: Guess a set of limbs $\mathcal{B}^{\prime}$ by calling LIMB (Section 4.1.1).
4: Guess a set of edges $\mathcal{R}^{\prime}$ by calling CONNECT (Section 4.1.2).
5: Find a set of edges $\mathcal{M}=\bigcup_{F\in\mathcal{F}}M_{F}$, where $M_{F}$ is a minimum matching in $G[Odd(G_{\mathcal{A}})\cap V(F)]$ and $G_{\mathcal{A}}=\mathcal{A}\cup\mathcal{B}^{\prime}\cup\mathcal{R}^{\prime}\cup\mathcal{F}$ for each $F\in\mathcal{F}$.
6: Get a TSP tour $T_{3}$ in $G$ by calling SHORTCUT (Section 4.1.3).
(a) The $\mathcal{A}$ and $\mathcal{B}$
(b) The guessed limbs in $\mathcal{B}^{\prime}$
(c) The chains in $\mathcal{R}$
(d) The minimal chains in $\mathcal{R}^{*}$
(e) The guessed edges in $\mathcal{R}^{\prime}$
(f) The edges in $\mathcal{M}$
Figure 3: An illustration of our ALG.3, where each black (resp., white) node represents a bad (resp., internal) vertex, each green node represents an anchor, and each ellipse represents the vertex set of a tree $F\in\mathcal{F}$. Moreover, each ellipse w.r.t. $F\in\mathcal{F}$ is divided into two parts, and the above part represents the potential vertex set $V_{F}$. In (a), the black (resp., red) edges represent the edges in $E(\mathcal{A})$ (resp., the limbs in $\mathcal{B}$); In (b), the red edges represent the guessed limbs in $\mathcal{B}^{\prime}$, and the white and green nodes incident to red edges represent guessed anchors; In (c), the blue edges represent the edges in $E(\mathcal{R})$; In (d), the blue edges represent the edges in $E(\mathcal{R}^{*})$, where $\mathcal{R}^{*}$ is a set of minimal chains obtained by shortcutting the internal vertices of the chains in $\mathcal{R}$; In (e), the blue edges represent the guessed edges in $\mathcal{R}^{\prime}$; In (f), the purple edges represent the edges in $\mathcal{M}$.
We remark that instead of guessing a set of limbs $\mathcal{B}^{\prime}$, the previous FPT $11$-approximation algorithm Zhou et al. (2024) mainly uses the spanning $|\mathcal{A}|$-forest $\mathcal{F}$ to guess a set of edges $E^{\prime}$ such that $\mathcal{A}\cup E^{\prime}$ forms a TSP tour in $G[V_{b}\cup V(E^{\prime})]$. Then, it further finds a special spanning forest $\mathcal{F}^{\prime}$ and then obtains a TSP tour by taking shortcuts on $\mathcal{A}\cup E^{\prime}\cup 2\mathcal{F}^{\prime}$, which is similar to the FPT $3$-approximation algorithm for TSP parameterized by $p$, as mentioned in the Section 3.2.
Next, we show the sub-algorithms LIMB and CONNECT in Steps 3 and 4, the property of $\mathcal{M}$ in Step 5 and the sub-algorithm SHORTCUT in Step 6, respectively.
4.1.1 The sub-algorithm: LIMB
Given $\mathcal{A}$ and $\mathcal{F}$, we assume that $T^{*}=\overline{a_{1}...b_{1}...a_{|\mathcal{A}|}...b_{|\mathcal{A}|}...}$, where $a_{i}...b_{i}\in\mathcal{A}$ for each $i$. Denote the number of anchors between $b_{i}$ and $a_{i+1}$ in $T^{*}$ by $f_{i}$, where $f_{i}\in\{1,2\}$. For each tree $F\in\mathcal{F}$, initialize a potential (vertex) set $V_{F}\coloneqq\emptyset$.
The main idea of guessing $\mathcal{B}^{\prime}$ is as follows. First, guess the values of $f_{1},...,f_{|\mathcal{A}|}$. By using these values, we know the positions of anchors based on $\mathcal{A}$. Then, we use “for each anchor” to express that we consider each anchor w.r.t. its position in the order it appears in $\overline{a_{1}...b_{1}...a_{|\mathcal{A}|}...b_{|\mathcal{A}|}...}$, respectively. Therefore, for each anchor, we know whether it is a single or pair anchor, and we also know the bad vertices incident to it in $T^{*}$. For example, we consider the $T^{*}$ in Fig. 1. If we guess correctly that $f_{1}=2$, $f_{2}=1$, and $f_{3}=2$, the anchors are pair, pair, single, pair, and pair, respectively, and the set of bad vertices incident to anchors are $\{x_{4}\}$, $\{x_{5}\}$, $\{x_{6},x_{7}\}$, $\{x_{7}\}$, $\{x_{1}\}$, respectively.
Next, for each anchor $x$, guess $F_{x}\in\mathcal{F}$ with $x\in V(F)$, carefully select a set of potential vertices $V_{x}\subseteq V(F_{x})$ and set $V_{F_{x}}\coloneqq V_{F_{x}}\cup V_{x}$. We will show that the size of each final potential vertex set is $\mathcal{O}(q^{2})$, and then, by guessing anchors in these potential vertex sets, we can obtain a set of limbs $\mathcal{B}^{\prime}$ satisfying some good properties. See details in LIMB.
Algorithm 4 LIMB
Input: An instance $G=(V=V_{g}\cup V_{b},E,w)$, $\mathcal{A}$, and $\mathcal{F}$.
Output: A set of limbs $\mathcal{B}^{\prime}$.
1: Initialize $V_{F}\coloneqq\emptyset$ for each $F\in\mathcal{F}$, and $\mathcal{B}^{\prime}=\emptyset$;
2: Guess the values of $f_{1},...,f_{|\mathcal{A}|}$.
3:for each anchor $x\in V_{a}$ do
4:     Guess the unique tree $F_{x}\in\mathcal{F}$ with $x\in V(F_{x})$, and let $n_{x}=\min\{2q,|V(F_{x})|\}$.
5:    if $x$ is a single anchor then
6:          Let $y$ and $z$ be the bad vertices incident to $x$ in $T^{*}$.
7:          Obtain a set of $n_{x}$ vertices $V_{x}$ by selecting $x^{\prime}\in V(F_{x})$ in the increasing order of $w(x^{\prime},y)+w(x^{\prime},z)$.
8:    else
9:          Let $y$ be the unique bad vertex incident to $x$ in $T^{*}$.
10:          Obtain a set of $n_{x}$ vertices $V_{x}$ by selecting $x^{\prime}\in V(F_{x})$ in the increasing order of $w(x^{\prime},y)$.
11:    end if
12:     Update $V_{F_{x}}\coloneqq V_{F_{x}}\cup V_{x}$.
13:end for
14:for each anchor $x\in V_{a}$ do
15:     Guess $x$ by using a vertex $x^{\prime}\in V_{F_{x}}$.
16:    if $x$ is a single anchor then
17:          Let $y$ and $z$ be the bad vertices incident to $x$ in $T^{*}$.
18:          Update $\mathcal{B}^{\prime}\coloneqq\mathcal{B}^{\prime}\cup\{x^{\prime}y,x^{\prime}z\}$$\triangleright$ $x^{\prime}y$ and $x^{\prime}z$ are the limbs w.r.t. $x^{\prime}$.
19:    else
20:          Let $y$ be the unique bad vertices incident to $x$ in $T^{*}$.
21:          Update $\mathcal{B}^{\prime}\coloneqq\mathcal{B}^{\prime}\cup\{x^{\prime}y\}$$\triangleright$ $x^{\prime}y$ is the limb w.r.t. $x^{\prime}$.
22:    end if
23:end for
Lemma 8 ($\clubsuit$).
There exists one configuration for the guessed anchors such that
  1. 1.
    the guessed anchors are pair-wise different;
  2. 2.
    for each anchor $x\in V_{a}$, the guessed anchor $x^{\prime}$ for $x$ satisfies $x^{\prime}=x$ if $x\in V_{F_{x}}$, and $x^{\prime}\in V_{x}\subseteq V_{F_{x}}$ otherwise;
  3. 3.
    the set of all limbs $\mathcal{B}^{\prime}$ w.r.t. guessed anchors satisfies $w(\mathcal{B}^{\prime})\leq w(\mathcal{B})$.
Moreover, LIMB can compute such a configuration in $\mathcal{O}(2^{q}\cdot q^{2q}\cdot(4q^{2})^{2q})$ loops, where each loop takes $n^{\mathcal{O}(1)}$ time.
4.1.2 The sub-algorithm: CONNECT
The set $\mathcal{A}\cup\mathcal{B}^{\prime}\cup\mathcal{F}$ may form a disconnected graph. For example, each tree $F\in\mathcal{F}$ with $V(F)\cap V_{a}=\emptyset$ must form a component (see (b) in Fig.3). We introduce the sub-algorithm CONNECT, which augments $\mathcal{A}\cup\mathcal{B}^{\prime}\cup\mathcal{F}$ into a connected graph $\mathcal{A}\cup\mathcal{B}^{\prime}\cup\mathcal{R}^{\prime}\cup\mathcal{F}$ by guessing a set of edges $\mathcal{R}^{\prime}$.
The main idea of guessing $\mathcal{R}^{\prime}$ is as follows. It can be verified that augmenting $\mathcal{A}\cup\mathcal{B}^{\prime}\cup\mathcal{F}$ by adding the set of good chains $\mathcal{R}$ ensures a connected graph $\mathcal{A}\cup\mathcal{B}^{\prime}\cup\mathcal{R}\cup\mathcal{F}$. However, since $|V(\mathcal{R})|=|V_{g}|=n-q$, we may not be able to guess $\mathcal{R}$ in FPT time. To overcome this issue, we turn to guess the set of trees $\mathcal{F}_{R}\subseteq\mathcal{F}$ connected by $R$ for each $R\in\mathcal{R}$, as $|\mathcal{F}|\leq q$ by (3). Then, we obtain a complete graph $\widetilde{G}=(\widetilde{V},\widetilde{E},\widetilde{w})$ by contracting each tree $F\in\mathcal{F}$ into a vertex $u_{F}$ and setting the weight between $u_{F}$ and $u_{F^{\prime}}$ as $\widetilde{w}(u_{F},u_{F^{\prime}})=\min_{v\in V(F),v^{\prime}\in V(F^{\prime})}w(v,v^{\prime})$. We observe that for each chain $R\in\mathcal{R}$, by shortcutting the internal vertices, we can obtain a minimal chain $R^{*}$ connecting all trees in $\mathcal{F}_{R}$ with a non-increasing weight, where $\widetilde{V}_{R}=\{u_{F}\mid F\in\mathcal{F}_{R}\}$. Moreover, $R^{*}$ corresponds to a TSP path/tour in $\widetilde{G}[\widetilde{V}_{R}]$. Thus, we can use the DP method Bellman (1962); Held and Karp (1962) to compute an optimal TSP path/tour in $\widetilde{G}[\widetilde{V}_{R}]$ with a weight of at most $w(R^{*})\leq w(R)$, which corresponds to a set of edges in $G$, denoted by $E_{R}$, connecting all trees in $\mathcal{F}_{R}$ (see (e) in Fig.3). It holds that $w(E_{R})\leq w(R^{*})$. Therefore, we set $\mathcal{R}^{\prime}=\bigcup_{R\in\mathcal{R}}E_{R}$. The details are described in CONNECT.
Algorithm 5 CONNECT
Input: An instance $G=(V=V_{g}\cup V_{b},E,w)$, $\mathcal{A}$, $\mathcal{B}^{\prime}$, and $\mathcal{F}$.
Output: A set of edges $\mathcal{R}^{\prime}$.
1:Initialize $\mathcal{R}^{\prime}=\emptyset$.
2:Obtain a complete graph $\widetilde{G}=(\widetilde{V},\widetilde{E},\widetilde{w})$ by contracting each $F\in\mathcal{F}$ into a vertex $u_{F}$, where, for any $u_{F},u_{F^{\prime}}\in\widetilde{V}$, define $\widetilde{w}(u_{F},u_{F^{\prime}})=\min_{v\in V(F),v^{\prime}\in V(F^{\prime})}w(v,v^{\prime})$
3:for each good chain $R\in\mathcal{R}$ with $V(R)\geq 2$ do
4:    Denote the pair anchors w.r.t. $R$ by $x$ and $x^{\prime}$.
5:    Guess the set of trees $\mathcal{F}_{R}\subseteq\mathcal{F}$ connected by $R$, and let $\widetilde{V}_{R}=\{u_{F}\mid F\in\mathcal{F}_{R}\}$.
6:    Using the DP Bellman (1962); Held and Karp (1962), find a minimum TSP tour in $\widetilde{G}[\widetilde{V}_{R}]$ if $F_{x}=F_{x^{\prime}}$, and a minimum TSP path between $u_{F_{x}}$ and $u_{F_{x^{\prime}}}$ in $\widetilde{G}[\widetilde{V}_{R}]$ otherwise.
7:    Denote the set of edges in $G$ w.r.t. the TSP tour/path by $E_{R}$.
8:    Update $\mathcal{R}^{\prime}\coloneqq\mathcal{R}^{\prime}\cup E_{R}$.
9:end for
Lemma 9 ($\clubsuit$).
CONNECT can compute $\mathcal{R}^{\prime}$, a set of edges between good vertices, in $\mathcal{O}(2^{q^{2}})$ loops such that
  1. 1.
    $\mathcal{A}\cup\mathcal{B}^{\prime}\cup\mathcal{R}^{\prime}\cup\mathcal{F}$ forms a connected graph $G_{\mathcal{A}}$;
  2. 2.
    $|Odd(G_{\mathcal{A}})\cap V(F)|$ is even for each $F\in\mathcal{F}$;
  3. 3.
    $Odd(G_{\mathcal{A}})\cap V_{a}=\emptyset$;
  4. 4.
    $w(\mathcal{R}^{\prime})\leq w(\mathcal{R})$.
Moreover, each loop takes $\mathcal{O}(q^{2}\cdot 2^{q})\cdot n^{\mathcal{O}(1)}$ time.
4.1.3 The sub-algorithm: SHORTCUT
Given $G_{\mathcal{A}}=\mathcal{A}\cup\mathcal{B}^{\prime}\cup\mathcal{R}^{\prime}\cup\mathcal{F}$, Step 5 of ALG.3 obtains an Eulerian graph $G^{\prime}_{\mathcal{A}}=\mathcal{A}\cup\mathcal{B}^{\prime}\cup\mathcal{R}^{\prime}\cup\mathcal{F}\cup\mathcal{M}$ by using $\mathcal{M}$ to fix the parity of the degrees of the vertices in $Odd(G_{\mathcal{A}})$. By Lemma 9, this step is feasible.
To analyze the quality of $\mathcal{M}$, we use the following lemma.
Lemma 10 ($\clubsuit$).
It holds that $w(\mathcal{M})\leq w(\mathcal{F})$.
Since $G^{\prime}_{\mathcal{A}}=\mathcal{A}\cup\mathcal{B}^{\prime}\cup\mathcal{R}^{\prime}\cup\mathcal{F}\cup\mathcal{M}$ is an Eulerian graph, we use the sub-algorithm SHORTCUT to take shortcuts on $G^{\prime}_{\mathcal{A}}$ to obtain a TSP tour $T_{3}$ in $G$ with a non-increasing weight.
In $G^{\prime}_{\mathcal{A}}$, each guessed single anchor connects two bad chains. Since a triangle containing one bad vertex may violate the triangle inequality, we need to carefully shortcut each guessed single anchor. To achieve this, in SHORTCUT, we first obtain a new Eulerian graph $G^{\prime\prime}_{\mathcal{A}}$ by deleting some edges in $\mathcal{M}\cup\mathcal{F}$ and making a copy for some guessed single anchors. Then, we can use a similar idea in Lemma 6 to take shortcuts on $G^{\prime\prime}_{\mathcal{A}}$ to obtain a TSP tour $T_{3}$ with a non-increasing weight.
The graph $G^{\prime\prime}_{\mathcal{A}}$ is constructed as follows. For each $F\in\mathcal{F}$ with $V(F)$ containing one guessed single anchor, we consider the following two cases. If $V(F)$ consists only of guessed single anchors, we know that $V(F)$ contains no internal vertices and pair anchors, and thus no edges in $\mathcal{R}^{\prime}$ connect $F$. Then, we simply delete all edges in $E(F)\cup M_{F}$ which dose not affect the connectivity of the graph, and refer to each guessed single anchor as a marked bad vertex. Note that a marked bad vertex is a bad vertex. Otherwise, $V(F)$ contains at least one internal vertex or pair anchor, and thus there exists one edge in $\mathcal{R}^{\prime}$ or one limb in $\mathcal{B}^{\prime}$ w.r.t. the pair anchor that connects $F$. Then, for each guessed single anchor $x\in V(F)$, assume that the corresponding limbs are $xy$ and $xz$, where $y$ and $z$ are bad vertices. We make one copy $x^{\prime}$ for the guessed single anchor $x$, delete the edges $xy$ and $xz$, add the edges $x^{\prime}y$ and $x^{\prime}z$, and refer to $x^{\prime}$ as a marked bad vertex. An illustration of the above two cases can be found in Fig. 4. The details can be found in SHORTCUT.
(a) $V(F)$ consists only of guessed single anchors.
(b) $V(F)$ consists of guessed single anchors and at least one internal vertex or pair anchor.
Figure 4: An illustration of the two cases in SHORTCUT, where each black node represents a bad vertex, each green or white node represents a good vertex, each blue node represents a marked bad vertex, all black edges represent edges in $E(F)$, the purple edge represents the edge in $M_{F}$, and all red edges represent limbs. In (a), $F$ consists only of guessed single anchors, and the black and purple edges are all deleted; In (b), $x^{\prime}$ is the copy for $x$.
Algorithm 6 SHORTCUT
Input: $G=(V=V_{g}\cup V_{b},E,w)$, $\mathcal{A}$, $\mathcal{B}^{\prime}$, $\mathcal{R}^{\prime}$, $\mathcal{F}$, and $\mathcal{M}$.
Output: A TSP tour $T_{3}$ in $G$.
1:for each $F\in\mathcal{F}$ with $V(F)$ containing one guessed single anchor do
2:    if $V(F)$ consists only of guessed single anchors then
3:         Delete all edges in $E(F)\cup M_{F}$, and refer to all the guessed single anchors as marked bad vertices.
4:    else
5:         for each guessed single anchor $x\in V(F)$ do
6:             Denote the limbs w.r.t. $x$ by $xy$ and $xz$.
7:             Create a marked bad vertex $x^{\prime}$.
8:             Replace the edges $xy$ and $xz$ with $x^{\prime}y$ and $x^{\prime}z$.
9:         end for
10:    end if
11:end for
12:Denote the above graph by $G^{\prime\prime}_{\mathcal{A}}$.
13:Obtain a TSP tour $T_{3}$ in $G$ (see Lemma 12).
Lemma 11 ($\clubsuit$).
$G^{\prime\prime}_{\mathcal{A}}$ remains an Eulerian graph, where each bad vertex is a 2-degree vertex and each good vertex is incident to at most one bad vertex.
Lemma 12 ($\clubsuit$).
Given $G^{\prime\prime}_{\mathcal{A}}$, there is an $\mathcal{O}(n^{2})$-time algorithm to obtain a TSP tour $T_{3}$ in $G$ with weight at most $w(G^{\prime\prime}_{\mathcal{A}})$.
Theorem 13 ($\clubsuit$).
For TSP parameterized by $q$, ALG.3 has an FPT $3$-approximation ratio with runtime $2^{\mathcal{O}(q\log q)}\cdot n^{\mathcal{O}(1)}$.
5 Conclusion
In this paper, we introduce several novel techniques to develop improved FPT approximation algorithms for TSP parameterized by $p$ and $q$. An interesting direction for future work would be to extend these techniques to design FPT approximation algorithms for related problems parameterized by $p$ and $q$, such as asymmetric TSP Traub and Vygen (2022) and prize-collecting TSP Blauth and Nägele (2023).
When the parameter $p$ (or $q$) is a small integer close to zero, our FPT approximation ratio remains unchanged and is larger than the approximation ratio $\alpha$ of metric TSP. This naturally leads to the question: is it possible to design an $\alpha$-approximation algorithm, even with a running time of $n^{\mathcal{O}(p+1)}$ (or $n^{\mathcal{O}(q+1)}$)? Furthermore, for the parameter $q$, can we develop a constant approximation algorithm with a running time of only $2^{\mathcal{O}(q)}\cdot n^{\mathcal{O}(1)}$?
Appendix
Proof of Lemma 4
Lemma 4. $F_{\mathcal{A}}$ in ALG.2 is a minimum constrained spanning tree in $G$ such that $E(\mathcal{A})\subseteq E(F_{\mathcal{A}})$, each $i$-degree vertex in $\mathcal{A}$, where $i\in\{0,1,2\}$, is incident to exactly $i$ bad vertices in $F_{\mathcal{A}}$, and each $2$-degree vertex in $\mathcal{A}$ remains a $2$-degree vertex in $F_{\mathcal{A}}$. Moreover, it holds that $\mbox{OPT}\geq w(F_{\mathcal{A}})$.
Proof.
First, we show that $F_{\mathcal{A}}$ is a constrained spanning tree in $G$. By Step 2 of ALG.2, $F$ in $G$, which corresponds to $\widetilde{F}$ in $\widetilde{G}$, consists only of edges between good vertices and edges between one good vertex and one bad vertex in $V^{0}_{b}\cup V^{1}_{b}$. Moreover, since $\widetilde{F}$ is a spanning tree in $\widetilde{G}$, we know that $F_{\mathcal{A}}=F\cup\mathcal{A}$ is a constrained spanning tree in $G$.
Then, we show that $F_{\mathcal{A}}$ is a minimum constrained spanning tree in $G$. Consider any minimum constrained spanning tree $F^{\prime}$ in $G$. By definition, $E(\mathcal{A})\subseteq E(F^{\prime})$, and the set $E(F^{\prime})\setminus E(\mathcal{A})$ consists only of edges between good vertices and edges between one good vertex and one bad vertex in $V^{0}_{b}\cup V^{1}_{b}$. So, $E(F^{\prime})\setminus E(\mathcal{A})$ corresponds to a spanning tree in $\widetilde{G}$ with a non-increasing weight by Step 2 of ALG.2. Since $F_{\mathcal{A}}$ satisfies $w(F_{\mathcal{A}})=\widetilde{w}(\widetilde{F})+w(\mathcal{A})$ and $\widetilde{F}$ is a minimum spanning tree in $\widetilde{G}$, we know that $w(F_{\mathcal{A}})\leq w(F^{\prime})$.
Last, by arbitrarily deleting an edge incident to a good vertex in $E(T^{*})$, we obtain a constrained spanning tree in $G$ with weight at most OPT. Thus, we have $\mbox{OPT}\geq w(F_{\mathcal{A}})$. ∎
Proof of Lemma 5
Lemma 5. It holds that $\frac{1}{2}\cdot\mbox{OPT}\geq w(M_{\mathcal{A}})$.
Proof.
By Lemma 4, each $2$-degree vertex in $\mathcal{A}$ remains a $2$-degree vertex in $F_{\mathcal{A}}$. Thus, $Odd(F_{\mathcal{A}})\subseteq V^{0}_{b}\cup V^{1}_{b}\cup V_{g}$.
Next, we try to shortcut vertices not in $Odd(F_{\mathcal{A}})$ on the optimal TSP tour $T^{*}$ with a non-increasing weight. First, we consider vertices in $V_{b}$. For any bad chain $u...u^{\prime}\in\mathcal{A}$, if at least one terminal is not in $Odd(F_{\mathcal{A}})$, say $u$, there must exist a good vertex $v$ in $T^{*}$ that is incident to $u$, and then we can use $v$ to repeatedly shortcut all vertices on $u...u^{\prime}$ that are not in $Odd(F_{\mathcal{A}})$ without increasing the weight by Property 1; if both terminals are in $Odd(F_{\mathcal{A}})$, we do not shortcut the vertices on $u...u^{\prime}$. Then, we consider vertices in $V_{g}$. By Property 1, we can shortcut all vertices in $V_{g}$ that are not in $Odd(F_{\mathcal{A}})$ without increasing the weight.
Thus, we can obtain a simple tour $T$ with $\mbox{OPT}\geq w(T)$, which consists of all vertices in $Odd(F_{\mathcal{A}})$ as well as all bad vertices in $V^{2}_{b}$ that are part of bad chains where both terminals belong to $Odd(F_{\mathcal{A}})$. By Step 5 of ALG.2, $T$ corresponds to a TSP tour $T^{\prime}$ in $G^{\prime}$, where $w(T)=w^{\prime}(T^{\prime})$. Clearly, $|Odd(F_{\mathcal{A}})|$ is even. Therefore, $T^{\prime}$ can be decomposed into two edge-disjoint matchings in $G^{\prime}$. Since $M^{\prime}_{\mathcal{A}}$ is a minimum matching in $G^{\prime}$ and $w(M_{\mathcal{A}})=w(M^{\prime}_{\mathcal{A}})$ by Step 6 of ALG.2, we have $\frac{1}{2}\cdot\mbox{OPT}\geq w(M_{\mathcal{A}})$. ∎
Proof of Theorem 7
Theorem 7. For TSP parameterized by $p$, ALG.2 has an FPT $1.5$-approximation ratio with runtime $2^{\mathcal{O}(p\log p)}\cdot n^{\mathcal{O}(1)}$.
Proof.
For the solution quality, by Lemmas 4, 5, and 6, we know that ALG.2 achieves an approximation ratio of $1.5$.
For the running time, Step 1 takes $\mathcal{O}(2^{p}\cdot p!)\subseteq 2^{\mathcal{O}(p\log p)}$ loops Bampis et al. [2024], and Steps 2-7 take $\mathcal{O}(n^{3})$ time, dominated by computing the minimum matching Lawler [1976]. Thus, the overall running time is $2^{\mathcal{O}(p\log p)}\cdot n^{\mathcal{O}(1)}$. ∎
Proof of Lemma 8
Lemma 8. There exists one configuration for the guessed anchors such that
  1. 1.
    the guessed anchors are pair-wise different;
  2. 2.
    for each anchor $x\in V_{a}$, the guessed anchor $x^{\prime}$ for $x$ satisfies $x^{\prime}=x$ if $x\in V_{F_{x}}$, and $x^{\prime}\in V_{x}\subseteq V_{F_{x}}$ otherwise;
  3. 3.
    the set of all limbs $\mathcal{B}^{\prime}$ w.r.t. guessed anchors satisfies $w(\mathcal{B}^{\prime})\leq w(\mathcal{B})$.
Moreover, LIMB can compute such a configuration in $\mathcal{O}(2^{q}\cdot q^{2q}\cdot(4q^{2})^{2q})$ loops, where each loop takes $n^{\mathcal{O}(1)}$ time.
Proof.
First, assume that both Steps 2 and 4 guess correctly.
Next, we further create a configuration for the guessed anchors satisfying the three conditions in the lemma.
Let $V_{a}=V^{0}_{a}\cup V^{1}_{a}$, where $V^{0}_{a}=\{x\mid x\in V_{a},x\notin V_{x}\}$, and $V^{1}_{a}=\{x\mid x\in V_{a},x\in V_{x}\}$. For each anchor $x$, let $x^{\prime}$ denote the guessed anchor w.r.t. $x$. We show how to set $x^{\prime}$.
We first consider the anchors in $V^{1}_{a}$. For each $x\in V^{1}_{a}$, since $x\in V_{x}$, we let $x^{\prime}=x$. Clearly, the weight of limbs w.r.t. $x^{\prime}$ equals to the weight of limbs w.r.t. $x$.
Next, we consider the anchors in $V^{0}_{a}$. For each $x\in V^{0}_{a}$, since $x\notin V_{x}$ and $V_{x}\subseteq V(F_{x})$, we have $V(F_{x})\setminus V_{x}\neq\emptyset$ and thus $|V(F_{x})|\geq 2q$. By Steps 7 and 10, we have $|V_{x}|=2q$. Since $|V_{a}|\leq 2q$ by (2), at least one vertex $x^{\prime\prime}\in V_{x}$ is still unused for guessing anchors, and then we let $x^{\prime}=x^{\prime\prime}$. By Steps 7 and 10, the weight of limbs w.r.t. $x^{\prime}$ is at most the weight of limbs w.r.t. $x$.
The above configuration clearly satisfies the first two conditions in the lemma. For any pair of guessed anchors $x^{\prime}$ and $x^{\prime\prime}$, the set of limbs w.r.t. $x^{\prime}$ is disjoint with the set of limbs w.r.t. $x^{\prime\prime}$. Since $\mathcal{B}^{\prime}$ is the set of all limbs w.r.t. all guessed anchors and $\mathcal{B}$ is the set of all limbs w.r.t. all anchors in $V_{a}$, we have $w(\mathcal{B}^{\prime})\leq w(\mathcal{B})$.
Next, we analyze the number of loops in LIMB. Since $f_{i}\in\{1,2\}$ for each $i\leq|\mathcal{A}|$, and $|\mathcal{A}|\leq q$ by (2), Step 2 needs to use $\mathcal{O}(2^{q})$ loops to correctly guess the values of $f_{1},...,f_{|\mathcal{A}|}$. Since $|V_{a}|\leq 2q$ by Step 4 and $|\mathcal{F}|\leq q$ by (3), LIMB needs to further use $\mathcal{O}(q^{2q})$ loops correctly guess the unique tree $F_{x}\in\mathcal{F}$ with $x\in V(F_{x})$ for each anchor $x\in V_{a}$. Moreover, for each anchor $x\in V_{a}$, we have $|V_{x}|\leq 2q$ by Steps 7 and 10. Thus, for each $F\in\mathcal{F}$, we have $|V_{F}|\leq 4q^{2}$. Since LIMB guesses $x$ by using a vertex from $V_{F_{x}}$ for each anchor $x\in V_{a}$ by Step 15, there are further $\mathcal{O}((4q^{2})^{2q})$ possible outcomes for the guessed anchors.
Overall, there are at most $\mathcal{O}(2^{q}\cdot q^{2q}\cdot(4q^{2})^{2q})$ possible configurations for the guessed anchors. Thus, the desired configuration can be computed in $\mathcal{O}(2^{q}\cdot q^{2q}\cdot(4q^{2})^{2q})$ loops. Clearly, each loop takes $n^{\mathcal{O}(1)}$ time. ∎
Proof of Lemma 9
Lemma 9. CONNECT can compute $\mathcal{R}^{\prime}$, a set of edges between good vertices, in $\mathcal{O}(2^{q^{2}})$ loops such that
  1. 1.
    $\mathcal{A}\cup\mathcal{B}^{\prime}\cup\mathcal{R}^{\prime}\cup\mathcal{F}$ forms a connected graph $G_{\mathcal{A}}$;
  2. 2.
    $|Odd(G_{\mathcal{A}})\cap V(F)|$ is even for each $F\in\mathcal{F}$;
  3. 3.
    $Odd(G_{\mathcal{A}})\cap V_{a}=\emptyset$;
  4. 4.
    $w(\mathcal{R}^{\prime})\leq w(\mathcal{R})$.
Moreover, each loop takes $\mathcal{O}(q^{2}\cdot 2^{q})\cdot n^{\mathcal{O}(1)}$ time.
Proof.
Consider the number of loops in CONNECT, and the running time in each loop.
Since $|\mathcal{F}|\leq q$ by (3), each good chain $R\in\mathcal{R}$ connects at most $q$ trees in $\mathcal{F}$, i.e., $|\mathcal{F}_{R}|\leq q$. Moreover, since $|\mathcal{R}|\leq q$ by (2), to guess $\mathcal{F}_{R}$ correctly for each $R\in\mathcal{R}$, CONNECT needs to try $(2^{q})^{q}$ possibilities, and thus there exists $\mathcal{O}(2^{q^{2}})$ loops. Moreover, given $\mathcal{F}_{R}$ for each $R\in\mathcal{R}$, since CONNECT computes a TSP tour/path by using the DP method Bellman [1962]; Held and Karp [1962], this takes $\mathcal{O}(q^{2}\cdot 2^{q})$ time. Thus, each loop takes $\mathcal{O}(q^{2}\cdot 2^{q})\cdot n^{\mathcal{O}(1)}$ time.
Then, consider the properties of $\mathcal{R}^{\prime}$.
First, we prove that $\mathcal{A}\cup\mathcal{B}^{\prime}\cup\mathcal{R}^{\prime}\cup\mathcal{F}$ forms a connected graph. Since $T^{*}=\mathcal{A}\cup\mathcal{B}\cup\mathcal{R}$, we know that $\mathcal{A}\cup\mathcal{B}\cup\mathcal{R}$ is a connected graph. By Lemma 8, it holds that for any $vx\in\mathcal{B}$, where $v$ is a bad vertex and $x$ is an anchor, the guessed anchor $x^{\prime}$ for $x$ satisfies $x^{\prime}\in V_{x}\subseteq V_{F_{x}}$, and then there exists a limb $vx^{\prime}\in\mathcal{B}^{\prime}$ connecting $v$ and $F_{x}$. Thus, $\mathcal{A}\cup\mathcal{B}^{\prime}\cup\mathcal{R}\cup\mathcal{F}$ is also a connected graph. For each $R\in\mathcal{R}$ with $V(R)\geq 2$, it connects all trees in $\mathcal{F}_{R}$. By CONNECT, there exists a set of edges $E_{R}$ w.r.t. $R\in\mathcal{R}$ connecting all trees in $\mathcal{F}_{R}$, and thus $\mathcal{A}\cup\mathcal{B}^{\prime}\cup\mathcal{R}^{\prime}\cup\mathcal{F}$ forms a connected graph $G_{\mathcal{A}}$.
Then, to prove that $|Odd(G_{\mathcal{A}})\cap V(F)|$ is even for each $F\in\mathcal{F}$ and $Odd(G_{\mathcal{A}})\cap V_{a}=\emptyset$, it is sufficient to prove that each bad vertex in $G_{\mathcal{A}}$ is a 2-degree vertex, and $\mathcal{A}\cup\mathcal{B}^{\prime}\cup\mathcal{R}^{\prime}$ corresponds to an Eulerian graph, where for each $F\in\mathcal{F}$ all vertices in $V(F)$ are contacted into a single vertex. Since $\mathcal{R}^{\prime}$ and $\mathcal{F}$ only contain edges between good vertices, each bad vertex in $G_{\mathcal{A}}$ is a 2-degree vertex. Moreover, by CONNECT, it can be verified that $\mathcal{A}\cup\mathcal{B}^{\prime}\cup\mathcal{R}^{\prime}$ forms an Eulerian graph (see (e) in Fig. 3). Thus, these two conditions hold.
Last, for each $R\in\mathcal{R}$, it connects all trees in $\mathcal{F}_{R}$. Recall the minimal chain $R^{*}$, obtained by shortcutting the internal vertices in $R$, which satisfies $w(R^{*})\leq w(R)$. Since $R^{*}$ corresponds to a TSP tour/path in $\widetilde{G}[\widetilde{V}_{R}]$ and CONNECT uses the DP method to compute a minimum TSP tour/path, we know that $E_{R}$, the set of edges w.r.t. the TSP tour/path, has a weight of at most $w(E_{R})\leq w(R^{*})\leq w(R)$. Therefore, we have $w(\mathcal{R}^{\prime})\leq w(R)$. ∎
Proof of Lemma 10
We first prove the following claim.
Claim 14.
For any tree $T$ with any $S\subseteq V(T)$ and $|S|\bmod 2=0$, the minimum matching $M_{S}$ on $S$ satisfies that $w(M_{S})\leq w(T)$.
Proof.
Since $|S|\bmod 2=0$, by arbitrarily pairing all vertices in $S$, we obtain a set of paths $\mathcal{B}$, where each $v\sim v^{\prime}\in\mathcal{B}$ denotes the unique path between $v$ and $v^{\prime}$ on $T$. Assume that there exists two paths $v...xy...v^{\prime},u...xy...u^{\prime}\in\mathcal{B}$ sharing at least one edge $xy$. We can modify these two paths into the paths $v\sim u$ and $v^{\prime}\sim u^{\prime}$, and the total number of edges on these two paths is reduced by at least 1. Therefore, by repeatedly performing the above procedure, we can ensure that $\mathcal{B}$ is a set of edge-disjoint paths.
By the triangle inequality, we have $w(v,v^{\prime})\leq w(v\sim v^{\prime})$ for any $v\sim v^{\prime}\in\mathcal{B}$. Thus, there exists a matching $\{vv^{\prime}\mid v\sim v^{\prime}\in\mathcal{B}\}$ with weight at most $w(\mathcal{B})\leq w(T)$. Since $M_{S}$ is the minimum matching, we have $w(M_{S})\leq w(T)$. ∎
Then, we are ready to prove Lemma 10.
Lemma 10. It holds that $w(\mathcal{M})\leq w(\mathcal{F})$.
Proof.
By Step 5 of ALG.3, $\mathcal{M}$ is obtained by finding a minimum matching $M_{F}$ in $G[Odd(G_{\mathcal{A}})\cap V(F)]$ for each $F\in\mathcal{F}$. Therefore, Lemma 10 follows directly from Lemma 9 and Claim 14. ∎
Proof of Lemma 11
Lemma 11. $G^{\prime\prime}_{\mathcal{A}}$ remains an Eulerian graph, where each bad vertex is a 2-degree vertex and each good vertex is incident to at most one bad vertex.
Proof.
In $G^{\prime}_{\mathcal{A}}=\mathcal{A}\cup\mathcal{B}^{\prime}\cup\mathcal{R}^{\prime}\cup\mathcal{F}\cup\mathcal{M}$, it can be verified that each bad vertex is a 2-degree vertex and each good vertex is incident to at most two bad vertices. Moreover, each good vertex that is incident to two bad vertices must be a guessed single anchor. Therefore, after processing all guessed single anchors in SHORTCUT, each bad vertex is a 2-degree vertex and each good vertex is incident to at most one bad vertex in $G^{\prime\prime}_{\mathcal{A}}$. In fact, each good vertex that is incident to one bad vertex must be a pair anchor.
In $G^{\prime\prime}_{\mathcal{A}}$, each vertex clearly remains an even degree vertex. Moreover, by CONNECT, it can be verified that $G^{\prime\prime}_{\mathcal{A}}$ remains connected, and thus $G^{\prime\prime}_{\mathcal{A}}$ remains an Eulerian graph. ∎
Proof of Lemma 12
We first prove the following claim.
Claim 15.
Given any Eulerian graph $G^{\prime}$ with $\mathcal{O}(m)$ edges, where each bad vertex is a 2-degree vertex and each good vertex is incident to at most one bad vertex, there is an $\mathcal{O}(m)$-time algorithm to obtain a TSP tour $T^{\prime}$ with a weight of at most $w(G^{\prime})$. Moreover, each good vertex that is incident to no bad vertices in $G^{\prime}$ is still incident to no bad vertices in $T^{\prime}$.
Proof.
Since $G^{\prime}$ is an Eulerian graph and it consists of $\mathcal{O}(m)$ edges, an Eulerian tour $T$ can be computed in $\mathcal{O}(m)$ time Cormen et al. [2022].
Since each bad vertex is a 2-degree vertex, it appears only once in $T$. If a vertex $a$ appears multiple times in $T$, it must be a good vertex. Moreover, since it is incident to at most one bad vertex in $G^{\prime}$, every appearance of $a$, except for one, is incident to two good vertices in $T^{\prime}$, and each such appearance can be shortcut without increasing the weight. Thus, we can obtain a TSP tour $T^{\prime}$ in $G^{\prime}$ with a weight of at most $w(G^{\prime})$, and the running time is $\mathcal{O}(m)$.
Since we shortcut a vertex $a$ only if $a$ is incident to two good vertices, the set of edges between one good vertex and one bad vertex remain unchanged in $T^{\prime}$. Therefore, each good vertex that is incident to no bad vertices in $G^{\prime}$ is still incident to no bad vertices in $T^{\prime}$. ∎
Then, we are ready to prove Lemma 12.
Lemma 12. Given $G^{\prime\prime}_{\mathcal{A}}$, there is an $\mathcal{O}(n^{2})$-time algorithm to obtain a TSP tour $T_{3}$ in $G$ with weight at most $w(G^{\prime\prime}_{\mathcal{A}})$.
Proof.
It can be verified that $|E(\mathcal{A})|\leq n$, $|\mathcal{B}^{\prime}|\leq n$, $|\mathcal{R}^{\prime}|\leq q^{2}\leq n^{2}$, $|E(\mathcal{F})|\leq n$, and $|\mathcal{M}|\leq n$. Thus, $G^{\prime\prime}_{\mathcal{A}}$ contains at most $\mathcal{O}(n^{2})$ edges.
By Lemma 11 and Claim 15, we can take shortcuts on $G^{\prime\prime}_{\mathcal{A}}$ to obtain a TSP tour $T$ with $w(T)\leq w(G^{\prime\prime}_{\mathcal{A}})$ in $\mathcal{O}(n)$ time. Note that $T$ is not a TSP tour in $G$ since, for each guessed single anchor $x$ in Step 5, it contains both $x$ and its copy $x^{\prime}$. By SHORTCUT, $x$ is incident to only good vertices in $G^{\prime\prime}_{\mathcal{A}}$, and thus it is still incident to only good vertices in $T$ by Claim 15. Therefore, we can keep $x^{\prime}$ and shortcut $x$ without increasing the weight. Therefore, given $T$, we can further obtain a TSP tour $T_{3}$ with $w(T_{3})\leq w(T)\leq w(G^{\prime\prime}_{\mathcal{A}})$ in $G$.
The overall running time is still $\mathcal{O}(n^{2})$. ∎
Proof of Theorem 13
Theorem 13. For TSP parameterized by $q$, ALG.3 has an FPT $3$-approximation ratio with runtime $2^{\mathcal{O}(q\log q)}\cdot n^{\mathcal{O}(1)}$.
Proof.
First, we analyze the quality of $T_{3}$.
By SHORTCUT and Lemma 12, we know that
$w(T_{3})$ $\leq w(G^{\prime\prime}_{\mathcal{A}})$
$\leq w(G^{\prime}_{\mathcal{A}})$
$=w(\mathcal{A})+w(\mathcal{B}^{\prime})+w(\mathcal{R}^{\prime})+w(\mathcal{F})+w(\mathcal{M}).$
Then, by (1), (3), Lemmas 8 and 9, and Lemma 10, we have
$w(\mathcal{A})+w(\mathcal{B}^{\prime})+w(\mathcal{R}^{\prime})+w(\mathcal{F})+w(\mathcal{M})$
$\leq w(\mathcal{A})+w(\mathcal{B})+w(\mathcal{R})+w(\mathcal{F})+w(\mathcal{F})$
$\leq 3(w(\mathcal{A})+w(\mathcal{B})+w(\mathcal{R}))$
$=3\cdot\mbox{OPT}.$
Next, we analyze the running time.
Step 1 takes $\mathcal{O}(2^{q}\cdot q!)$ loops, where each loop takes $n^{\mathcal{O}(1)}$ time Bampis et al. [2024]. Step 2 takes $\mathcal{O}(n^{2}\log n)$ time Khachay and Neznakhina [2016]. By Lemma 8, Step 3 takes $\mathcal{O}(2^{q}\cdot q^{2q}\cdot(4q^{2})^{2q})$ loops, where each loop takes $n^{\mathcal{O}(1)}$ time. By Lemma 9, Step 4 takes $\mathcal{O}(2^{q^{2}})$ loops, where each loop takes $\mathcal{O}(q^{2}\cdot 2^{q})\cdot n^{\mathcal{O}(1)}$ time. Step 5 takes $\mathcal{O}(n^{3})$ time Lawler [1976]. By Lemma 12, Step 6 takes $\mathcal{O}(n^{2})$ time.
Therefore, there are $2^{\mathcal{O}(q\log q)}$ loops in total, where each loop takes at most $2^{\mathcal{O}(q)}\cdot n^{\mathcal{O}(1)}$ time. Thus, the overall running time is $2^{\mathcal{O}(q\log q)}\cdot 2^{\mathcal{O}(q)}\cdot n^{\mathcal{O}(1)}\subseteq 2^{\mathcal{O}(q\log q)}\cdot n^{\mathcal{O}(1)}$. ∎